(a) Let In = \int _0^1x^n\sqrt{1-x^2\:dx}, n ∈ N. Show that (n + 2)In = (n − 1)In−2, n ≥ 2.

(b) If g = sin(sin x), prove that \frac{d^2y}{dx^2}+tan\:x\frac{dy}{dx}+ycosx^2= 0.


1
Expert's answer
2021-05-25T17:46:07-0400

(a)


"I_n=\displaystyle\int_{0}^{1}x^n\sqrt{1-x^2}dx"

"\int x^n\sqrt{1-x^2}dx"

"\int udv=uv-\int vdu"

"u=x^{n-1}, du=(n-1)x^{n-2}dx"

"dv=x\sqrt{1-x^2}dx, v=\int x\sqrt{1-x^2}dx"

"=-\dfrac{1}{3}(1-x^2)^{{3 \over 2}}"

"\int x^n\sqrt{1-x^2}dx=x^{n-1}\big(-\dfrac{1}{3}(1-x^2)^{{3 \over 2}}\big)"

"+\dfrac{n-1}{3}\int x^{n-2}(1-x^2)\sqrt{1-x^2}dx"

"=x^{n-1}\big(-\dfrac{1}{3}(1-x^2)^{{3 \over 2}}\big)+\dfrac{n-1}{3}\int x^{n-2}\sqrt{1-x^2}dx"

"-\dfrac{n-1}{3}\int x^{n}\sqrt{1-x^2}dx"

"(n+2)\int x^n\sqrt{1-x^2}dx=x^{n-1}(1-x^2)^{{3 \over 2}}"

"+(n-1)\int x^{n-2}\sqrt{1-x^2}dx"

"(n+2)\displaystyle\int_{0}^{1}x^n\sqrt{1-x^2}dx"

"=\big[x^{n-1}(1-x^2)^{{3 \over 2}}\big]\begin{matrix}
 1 \\
 0
\end{matrix}+(n-1)\displaystyle\int_{0}^{1}x^{n-2}\sqrt{1-x^2}dx"

"(n+2)\displaystyle\int_{0}^{1}x^n\sqrt{1-x^2}dx"

"=(0-0)+(n-1)\displaystyle\int_{0}^{1}x^{n-2}\sqrt{1-x^2}dx"

Hence


"(n+2)I_n=(n-1)I_{n-2}, \ \ n\geq2"


(b)


"g=\sin(\sin x)"

"\dfrac{dg}{dx}=\cos(\sin x)\cdot\cos x"


"\dfrac{d^2 g}{dx^2 }=-\sin(\sin x)\cdot\cos^2 x-\cos(\sin x)\cdot\sin x"

"\dfrac{d^2 g}{dx^2 }+\tan x\cdot\dfrac{d g}{dx }+g\cdot \cos^2x"

"=-\sin(\sin x)\cdot\cos^2 x-\cos(\sin x)\cdot\sin x"

"+\tan x\cdot\cos(\sin x)\cdot\cos x"

"+\sin(\sin x)\cdot\cos^2x"

"=0"